OpenStudy (anonymous):

Where did I make the mistake? Suppose you prove the identity cot theta = csc theta/sec theta by working only on the write side of the equation. What would the final step look like? I only ask this because I had put; cot theta = 1/tan theta (since the reciprocal equals cot theta) but it was marked as inccorect. Would the final step just by cot theta = cot theta?

terenzreignz (terenzreignz):

You know, it helps when you reduce *everything* to sines and cosines I suggest you give that a try

OpenStudy (anonymous):

Okay, hold on

OpenStudy (anonymous):

I'm just a little confused because when I write the reciprocal identity for csc theta, it becomes a fraciton on top of another fraction. It that okay when solving for identities?

terenzreignz (terenzreignz):

Of course, whenever you get a fraction on top of the other, you can express them as one *simpler* fraction. Do you know how?

OpenStudy (anonymous):

I'm afraid not :(

terenzreignz (terenzreignz):

Well, that's why I'm here ;) When you get a fraction on top of another, say (a/b) / (c/d) you take the top fraction, and multiply it to the *reciprocal* of the bottom fraction (a/b) * (d/c) Can you take it from here and apply it to your problem? :)

OpenStudy (anonymous):

Right now, I have cot theta = (1/sin theta) / (1/cos theta), so I use the reciprocal of 1/cos theta and get theta? So it'd be cot theta = (1/sin theta) / cos theta? (:

terenzreignz (terenzreignz):

I thought you were only working on the right side?

OpenStudy (anonymous):

I ment * and get cos theta.

terenzreignz (terenzreignz):

with csc theta / sec theta

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Sorry Lol

terenzreignz (terenzreignz):

ok, so csc = 1/sin, right? (I'll no longer put the theta)

terenzreignz (terenzreignz):

what's sec?

OpenStudy (anonymous):

Yes, and alrighty

terenzreignz (terenzreignz):

You get it now? ;)

OpenStudy (anonymous):

sec is 1/cos

OpenStudy (anonymous):

A little more (:

terenzreignz (terenzreignz):

that's right, so... it boils down to (1/sin) / (1/cos) now apply what I showed you about complex fractions

OpenStudy (anonymous):

So if I apply what you showed about the complex fractions, would it look like this? (1/sin) / (cos) ? I'm sorry if It's taking me a minute to grasp this.

terenzreignz (terenzreignz):

ok, remember (a/b) / (c/d) = (a/b) * (d/c) just substitute ;)

OpenStudy (anonymous):

Ooohh, okay, didn't see the multiplication there (: Lol

terenzreignz (terenzreignz):

So, can you do it?

OpenStudy (anonymous):

Well, I have (1/sin) * cos but I'm not sure what follows...

OpenStudy (anonymous):

Would it become tangent?

terenzreignz (terenzreignz):

(1/sin) * (cos/1) = just multiply their numerators and denominators and you should come to an epiphany :D

OpenStudy (anonymous):

Ooooh, cotangnet? Because it'll be cos/sin, right?

OpenStudy (anonymous):

And that proves the identity, right? (:

terenzreignz (terenzreignz):

That's right Good job :)

OpenStudy (anonymous):

Thank you! I love how you made me think! Some people just give the answers and it doesn't help for future problems (: I appreciate your help!

terenzreignz (terenzreignz):

Thanks, I learned my lesson :) Uhh, it does help to try to reduce everything (if you can/want) to sines and cosines when proving identities

OpenStudy (anonymous):

That's what my textbook said, but I didn't know what it really ment until now (: Thanks again!

terenzreignz (terenzreignz):

No problem =D